Sales Mix and Break-Even Analysis

Jordan Company has fixed costs of $273,600. The unit selling price, variable cost per unit, and contribution margin per unit for the company's two products are provided below.

Product Selling Price Variable Cost per Unit Contribution Margin per Unit
Q $120 $70 $50
Z 80 70 10

The sales mix for products Q and Z is 50% and 50%, respectively. Determine the break-even point in units of Q and Z. If required, round your answers to the nearest whole number.

a. Product Q  units

b. Product Z  units
Answer #1
Weighted average unit Contribution Margin 30 =(50*50%)+(10*50%)
Overall break-even point 9120 =273600/30
a
Product Q units 4560 =9120*50%
b
Product Z units 4560 =9120*50%